An enterprising student has set up an internship clearinghouse for business students. Each student who uses the service fills out a form and lists up to 23 companies that he or she would like to have contacted. The clearinghouse has a choice of two methods to use for processing the forms. The traditional method requires about 16 minutes to review the form and arrange the information in the proper order for processing. Once this setup is done, it takes only three minutes per company requested to complete the processing. The other alternative uses an optical scan/retrieve system, which takes only two minutes to prepare but requires three minutes per company for completing the processing.

a. Calculate the process time under each method?

b. If it costs about the same amount per minute for processing with either of the two methods, which method is better?
Answer #1

Method 1:

For 1 company, 16 minutes to review and 3 minutes for processing,

So, M1 = 16R + 3P = 19 minutes,

for 23 companies, it takes 16 min to review and 3*23 = 69 minutes for processing,

So, M1 = 16R + 69P = 85 minutes

Method 2:

For 1 company, it takes 2 min to review and 3 minutes for processing,

so, M2 = 2R + 3P = 5 minutes

For 23 companies, it takes 2 min to review and 3*23 = 69 minutes for processing,

so, M2 = 2R + 69P = 71 minutes

So, It's better to use method 2 for any number of companies since method 2 resulted in lower times.

****Please please please LIKE THIS ANSWER, so that I can get a small benefit, Please****